# Regularization methods for solving hierarchical variational inequalities with complexity guarantees

### v1

@@ 2.2 Restarting Procedure
Given an anchor point $w$, Algorithm 1 produces a sequence of points $(v_{k})$ whose last iterate corresponds to the output of the function $\mathtt{KM}(v,\mathsf{T}^{(w,\beta)}_{\bullet},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$ and which provides a good candidate for a new anchor point, given the tracking property (2.2). Repeating this over time for a fixed number of iterations is the key updating step constituting our outer loop procedure. The mechanics of this procedure gives rise to an diagonal equilibrium tracking method (DANTE), formally described in Algorithm 2.
@@
$w_{0}\in\mathcal{H}$ and $(\tau_{0,k}),(\theta_{0,k})_{k}$ given.
@@
   Set $v=\mathtt{KM}(w_{n},\mathsf{T}^{(w_{n},\beta_{n})}_{n,\bullet},(\tau_{n,k})_{k},(\theta_{n,k})_{k},\varepsilon_{n})$.
@@
Note that DANTE is a method rather than an algorithm as we do not specify how the parameters required by the fixed-point iteration procedure $\mathtt{KM}(v,\mathsf{T}^{(w,\beta)}_{\bullet},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$ are set. This is a part of the user-defined input and should be decided based on the eventually known problem structure.
@@ 3.1 Convergence Analysis of Inner Loop
[Cortild:2024aa, Hypothesis 2.1] Assume $(\theta_{k})_{k}\subset(0,\overline{\theta}]$ and that $(\tau_{k})_{k}\subset[0,\overline{\tau}]$ is monotonically non-decreasing, where $\overline{\theta},\overline{\tau}\in(0,1)$. Moreover, assume that, for all $k\in\mathbb{N}$,
@@
where $Q_{k}=1-\theta_{k}(1-q_{k}^{2})$, and that $\overline{Q}\triangleq\sup_{k\in\mathbb{N}}Q_{k}<1$.
Under the above assumption, we can guarantee strong convergence of the iterates.
@@
The result follows by the square-summability of $(v_{k}-p)$, and the rate provided in Theorem 3.1. $\blacksquare$
@@
We assume that $\operatorname{dom}(\mathsf{T}_{k})\subset\operatorname{dom}(\mathsf{M})$ for all $k\in\mathbb{N}$.
@@
Assumption 4 might seem restrictive, but in fact, they are satisfied by many operators. For instance, if $\mathsf{M}=\mathsf{F}+\operatorname{\mathsf{NC}}_{\mathcal{K}}$, where $\operatorname{\mathsf{NC}}_{\mathcal{K}}$ is the normal cone of ${\mathcal{K}}=\operatorname{dom}(\mathsf{M})$ and $\mathsf{F}$ is a Lipschitz monotone operator, the forward-backward operator (FB) satisfies the desired conditions, as the resolvent of $\operatorname{\mathsf{NC}}_{\mathcal{K}}$ reduces to a projection onto $\mathcal{K}$}, such that $\mathsf{T}(x)\in{\mathcal{K}}$ for all $x\in\mathcal{H}$. The forward-backward operator and other examples are discussed in Section 4.
@@
Assume Assumptions 2, 3, and 4 hold. Then
@@
$C\triangleq\frac{2(1+\overline{\tau})}{\sqrt{\overline{Q}(1-\overline{\tau})(1-\overline{\theta})}}\cdot\operatorname{diam}(\operatorname{dom}(\mathsf{M})).$ (Def-$C$)
@@
$\mathbb{K}(\varepsilon)\leq\overline{K}_{\varepsilon}\triangleq\left\lceil 2\log(C/\varepsilon)/\log(1/\overline{Q})\right\rceil.$
@@
By [Cortild:2024aa, Theorem 2.1], we also know that weak convergence of the iterates $(v_{k})$ and the velocities $z_{k}$ is guaranteed when $\mathsf{T}_{k}$ is just nonexpansive. Moreover, if $\mathsf{T}_{k}$ is a fixed-point encoding of Problem (Aux), which has a unique solution $p$ by construction, we know that the sequence convergences weakly to this unique fixed-point $p$. Specifically, in the finite-dimensional setting, we know that $\|v_{k}-p\|\to 0$ and $\|v_{k+1}-z_{k}\|\to 0$, meaning that Algorithm 2 still can be applied. It is however important to note that Lemma 2.1 no longer holds, and that we do not have an appropriate stopping criterion to guarantee $\|w^{+}-\overline{u}_{\beta}(w)\|\leq\mathtt{e}$. Within our final numerical implementation in Section 5.3, we allow for nonexpansive operators, corresponding to $q=1$, heavily noting that the theoretical guarantees are weakened in this setting.
@@
Rather than assuming exact access to the operators $\mathsf{T}_{k}$, we could assume access to $\delta_{k}$-perturbations $\tilde{\mathsf{T}}_{k}$, with bounded inexactness model $0\leq\delta_{k}\leq\delta_{\max}$. Such an error structure is motivated by the fact that the exact map might in general not be available, or its computation may be very demanding. This occurs for instance when applying proximal methods to image deblurring with total variation [Chambolle:2004aa], or to structured sparsity regularization problems in machine learning and inverse problems [Zhao:2009aa]. In those cases, the proximity operator is usually computed using ad hoc algorithms, and therefore inexactly. The implications of inexact implementations are the following: Lemma 2.1 would still hold, by replacing $\varepsilon$ by $\max(\varepsilon,\delta_{\max})$, provided $\delta_{\max}$ is finite. Additionally, Theorem 3.1 would still hold [Cortild:2024aa], provided $(\theta_{k}\delta_{k})\in\ell^{2}$, but without an explicit rate. As such, Corollary 3.2 still holds, but again without rate guarantees. Specifically, we have no apriori upper bound on the required number of iterations required for Algorithm 1 to terminate, but we know it will terminate.
@@ 3.2 Convergence Analysis of Restarting Procedure
In the following, we assume that the operator $\mathsf{G}$ is $\mu$-strongly monotone with $\mu\geq 0$. The case in which $\mathsf{G}$ is merely monotone is recovered by setting $\mu=0$ throughout the subsequent analysis.
We introduce two gap functions, one associated with optimality and one with feasibility.
@@
$\operatorname{Gap}_{\operatorname{opt}}(u)\triangleq\sup_{v\in\mathcal{S}_{0}}\langle\mathsf{G}v,u-v\rangle.$ ($\operatorname{Gap}_{\operatorname{opt}}$)
We observe that if $u$ is feasible, namely if $u\in\mathcal{S}_{0}=\operatorname{Zer}(\mathsf{M})$, then the condition $\operatorname{Gap}_{\operatorname{opt}}(u)\leq 0$ is equivalent to $u\in\operatorname{Zer}(\mathsf{G}+\operatorname{\mathsf{NC}}_{\mathcal{S}_{0}})$, such that $u$ solves Problem (P). It is important to note, however, that $\operatorname{Gap}_{\operatorname{opt}}(u)\leq 0$ provides limited information when $u\notin\operatorname{Zer}(\mathsf{M})$. It is thus important to obtain a lower bound on the optimality gap.
The feasibility gap is motivated by the observation that, by maximal monotonicity of $\mathsf{M}$, for any $x\in\mathcal{H}$,
@@
As such, we define the feasibility gap $\operatorname{Gap}_{\operatorname{feas}}:\mathcal{H}\to\mathbb{R}$ as
$\operatorname{Gap}_{\operatorname{feas}}(x)=\sup_{y\in\operatorname{dom}(\mathsf{M}),v\in\mathsf{M}y}\langle v,x-y\rangle.$ ($\operatorname{Gap}_{\operatorname{feas}}$)
We note that $\operatorname{Gap}_{\operatorname{feas}}(x)\leq 0$ implies $x\in\operatorname{Zer}(\mathsf{M})$, and that $\operatorname{Gap}_{\operatorname{feas}}(x)\geq 0$ holds for all $x\in\mathcal{H}$ since, by assumption, the problem admits at least one solution; in particular, $\operatorname{Zer}(\mathsf{M})\neq\varnothing$. Since $\operatorname{Gap}_{\operatorname{opt}}$ and $\operatorname{Gap}_{\operatorname{feas}}$ are pointwise suprema of affine functions, both are weakly lower semicontinuous.
@@
The notion of weak sharpness of zero sets is closely related to the classical concept of weak sharpness for variational inequalities [marcotte_weak_1998]. Geometric conditions ensuring the error-bound characterization (3.2) can be found in [Al-Homidan:2017aa]. The recent papers [samadi_improved_2025] and [Alvarez:2001aa] rely also heavily on this assumption. An important class of examples arises when $\mathsf{M}$ is single-valued and $\rho=1$, notably in monotone linear complementarity problems in finite dimensions under nondegeneracy conditions [burke_weak_1993, pang1997error]. In the case where the lower-level problem is determined by a convex optimization problem, weak-sharpness implies an Hölderian errror bound, an assumption already imposed by [Cabot2005] in the context of hierarchical minimization. Specifically, let us assume that $\mathsf{M}=\partial f$ for a proper convex and lower semi-continuous function $f:\mathcal{H}\to\mathbb{R}\cup\{+\infty\}$. Then,
@@
$D_{\mathsf{M}}=\operatorname{diam}(\operatorname{dom}(\mathsf{M})),\hskip 17.00024pt\hskip 17.00024pt$ $C_{\mathsf{M}}=\sup_{x\in\operatorname{dom}(\mathsf{M})}\sup_{v\in\mathsf{M}x}\|v\|,\qquad$ $C_{\mathsf{G}}=\sup_{x\in\operatorname{dom}(\mathsf{M})}\mathsf{G}x.$
@@
We assume the following on the parameter sequences:
@@
For any $x\in\operatorname{dom}(\mathsf{M})$ and $v\in\mathsf{M}x$, it holds that
$\lambda_{n}\langle v,w_{n+1}-x\rangle+\lambda_{n}\beta_{n}\langle\mathsf{G}x,w_{n+1}-x\rangle\leq-\frac{\alpha\lambda_{n+1}}{2}\|w_{n+1}-x\|^{2}+\frac{\alpha\lambda_{n}}{2}\|w_{n}-x\|^{2}+C_{1}\lambda_{n}\mathtt{e}_{n},$ (3.4)
@@
As $\beta_{n}\mathsf{G}\overline{u}_{n+1}-\alpha(\overline{u}_{n+1}-w_{n})\in\mathsf{M}\overline{u}_{n+1}$ and $v\in\mathsf{M}x$, it follows by monotonicity of $\mathsf{M}$ and by $\mu$-monotonicity of $\mathsf{G}$ that
$\langle w_{n}-\overline{u}_{n+1},x-\overline{u}_{n+1}\rangle$ $\leq\frac{1}{\alpha}\langle v,x-\overline{u}_{n+1}\rangle+\frac{\beta_{n}}{\alpha}\langle\mathsf{G}\overline{u}_{n+1},x-\overline{u}_{n+1}\rangle$
$\leq\frac{1}{\alpha}\langle v,x-\overline{u}_{n+1}\rangle+\frac{\beta_{n}}{\alpha}\langle\mathsf{G}x,x-\overline{u}_{n+1}\rangle-\frac{\mu\beta_{n}}{\alpha}\|x-\overline{u}_{n+1}\|^{2}.$
@@
$\frac{1}{2}\|w_{n+1}-x\|^{2}-\frac{1}{2}\|w_{n}-x\|^{2}$ $\leq\frac{1}{\alpha}\langle v,x-\overline{u}_{n+1}\rangle+\frac{\beta_{n}}{\alpha}\langle\mathsf{G}x,x-\overline{u}_{n+1}\rangle-\frac{\mu\beta_{n}}{\alpha}\|{x-w_{n+1}}\|^{2}$
@@
$\left(\frac{1}{2}+\frac{\mu\beta_{n}}{\alpha}\right)\|w_{n+1}-x\|^{2}-\frac{1}{2}\|w_{n}-x\|^{2}\leq\frac{1}{\alpha}\langle v,x-\overline{u}_{n+1}\rangle+\frac{\beta_{n}}{\alpha}\langle\mathsf{G}x,x-\overline{u}_{n+1}\rangle+2\left(1+\frac{\mu\beta_{\max}}{\alpha}\right)(D_{\mathsf{M}}+\mathtt{e}_{\max})\mathtt{e}_{n}.$
@@
$\langle\mathsf{G}x,x-\overline{u}_{n+1}\rangle=\langle\mathsf{G}x,x-w_{n+1}\rangle+\langle\mathsf{G}x,w_{n+1}-\overline{u}_{n+1}\rangle\leq\langle\mathsf{G}x,x-w_{n+1}\rangle+C_{\mathsf{G}}\mathtt{e}_{n},$
@@
$\lambda_{n}\langle v,w_{n+1}-x\rangle$ $+\lambda_{n}\beta_{n}\langle\mathsf{G}x,w_{n+1}-x\rangle$
@@
$\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}\langle\mathsf{G}x,w_{n+1}-x\rangle$ $\leq\frac{-\alpha\lambda_{N}}{2}\|w_{N}-x\|^{2}+\frac{\alpha\lambda_{0}}{2}\|w_{0}-x\|^{2}+C_{1}\cdot\sum_{n=0}^{N-1}\lambda_{n}\mathtt{e}_{n}$
@@
$\langle\mathsf{G}x,\overline{w}_{N}-x\rangle\leq C_{2}\cdot\frac{1}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}+C_{1}\cdot\frac{\sum_{n=0}^{N-1}\lambda_{n}\mathtt{e}_{n}}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}.$
@@
$\langle\mathsf{G}x,w_{n+1}-x\rangle\geq-\|{\mathsf{G}x}\|\cdot\|{w_{n+1}-x}\|\geq-\|\mathsf{G}x\|\cdot(\|w_{n+1}-\overline{u}_{n+1}\|+\|\overline{u}_{n+1}-x\|)\geq-C_{\mathsf{G}}(\mathtt{e}_{n}+D_{\mathsf{M}}).$
It then follows $\langle\mathsf{G}x,w_{n+1}-x\rangle\leq C_{\mathsf{G}}(\mathtt{e}_{\max}+D_{\mathsf{M}})$. Summing for $n=0,\ldots,N-1$ then yields
@@
We may now combine the two previous lemmas to state our main theorem.
@@
    The upper bound is just Lemma 3.6. For the lower bound, we invoke the Cauchy-Schwarz inequality $\langle\mathsf{G}x,\overline{w}_{N}-x\rangle\geq\|{\mathsf{G}x}\|\cdot\|{\overline{w}_{N}-x}\|$, implying
@@
Let $(w_{n})$ be generated through Algorithm 2 when $\mathsf{G}$ is merely monotone (i.e. $\mu=0$). Let $\beta_{n}=(n+1)^{-b}$ for $b\in(0,1)$ and choose $\varepsilon_{n}$ such that $\mathtt{e}_{n}=o(\beta_{n})$. Then, for all $N\geq 2^{1/(1-b)}$, it holds that
@@
If, moreover, $\operatorname{Zer}(\mathsf{M})$ is weakly sharp with constant $\kappa$ of order $\rho\geq 1$, then it also holds that $\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{N})\geq\mathcal{O}(\operatorname{Gap}_{\operatorname{feas}}(\overline{w}_{N})^{1/\rho}).$
@@
3.3 Combined Complexity Analysis
We now finally perform the complete complexity analysis of our scheme DANTE by combining our bounds derived for Algorithms 1 and 2.
Before stating the final convergence result, we make one final assumption.

### v2

@@ 2.2 Restarting Procedure
Given an anchor point $w$, Algorithm 1 produces a sequence of points $(v_{k})_{k}$, whose last iterate corresponds to the output of the function $\mathtt{KM}(v,(\mathsf{T}^{(w,\beta)}_{k})_{k},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$. In view of the tracking property established in (2.3), we will use this point as the new anchor point for the inner-loop KM iteration. Repeating this over time for a fixed number of restarts with updated parameters required by the inner-loop, defines our outer loop procedure. The mechanics of this procedure gives rise to a diagonal equilibrium tracking method (DANTE), formally described in Algorithm 2.
@@
$w_{0}\in\mathcal{H}$ given initial point. $(\tau_{0,k}),(\theta_{0,k})_{k}$ given parameters and $N\geq 1$ given number of restarts.
@@
   Set $v=\mathtt{KM}(w_{n},(\mathsf{T}^{(w_{n},\beta_{n})}_{n,k})_{k},(\tau_{n,k})_{k},(\theta_{n,k})_{k},\varepsilon_{n})$.
@@
Note that DANTE is a method rather than a bona-fide algorithm as we do not specify how the parameters required by the fixed-point iteration $\mathtt{KM}(v,(\mathsf{T}^{(w,\beta)}_{k})_{k},(\tau_{k})_{k},(\theta_{k})_{k},\varepsilon)$ are updated. This is a part of the user-defined input and should be decided based on the eventually known problem structure.
@@ 3.1 Convergence Analysis of Inner Loop
[Cortild:2024aa, Hypothesis 2.1] Assume $(\theta_{k})_{k}\subset(0,\overline{\theta}]$ and that $(\tau_{k})_{k}\subset[0,\overline{\tau}]$ is monotonically non-decreasing, where $\overline{\theta},\overline{\tau}\in(0,1)$. Moreover,
@@
for all $k\in\mathbb{N}$, where $Q_{k}=1-\theta_{k}(1-q_{k}^{2})$, and that $\overline{Q}\triangleq\sup_{k\in\mathbb{N}}Q_{k}<1$.
Under the above assumption, we can guarantee strong convergence of the iterates at a linear rate.
@@
The result follows by the square-summability of $(v_{k}-p)_{k}$, and the rate provided in Theorem 3.1. $\blacksquare$
@@
We assume that $\operatorname{dom}(\mathsf{T}^{(w,\beta)}_{k})\subset\operatorname{dom}(\mathsf{M})$ for all $k\in\mathbb{N}$ and all $(w,\beta)$.
@@
Assumption 4 might seem restrictive, but in fact, they are satisfied by many operators. As a concrete example, assume that we know that the lower level is formulated in terms of the operator $\mathsf{M}=\mathsf{F}+\operatorname{\mathsf{NC}}_{\mathcal{K}}$, where $\operatorname{\mathsf{NC}}_{\mathcal{K}}$ is the normal cone of the closed convex set ${\mathcal{K}}=\operatorname{dom}(\mathsf{M})$ and $\mathsf{F}:\mathcal{H}\to\mathcal{H}$ is a Lipschitz continuous monotone operator. With this problem structure, a powerful fixed-point encoding strategy is the forward-backward operator. Adapted to the auxiliary problem (Aux), this method is defined in terms of the mapping
$\mathsf{T}^{(w,\beta)}_{k}(x)=\Pi_{\mathcal{K}}(x-\gamma_{k}(\mathsf{F}(x)+\beta\mathsf{G}(x)+\alpha(x-w))),$
$\gamma_{k}>0$ being a suitably chosen step-size parameter. Clearly, $\mathsf{T}^{(w,\beta)}_{k}(x)\in{\mathcal{K}}$ for all $x\in\mathcal{H}$. The forward-backward operator and other examples are discussed in detail in Section 4.
@@
Let Assumptions 2-4 be in place. Then,
@@
$C\triangleq\frac{2(1+\overline{\tau})}{\sqrt{\overline{Q}(1-\overline{\tau})(1-\overline{\theta})}}\cdot\operatorname{diam}(\operatorname{dom}(\mathsf{M})).$ (3.2)
@@
$\mathbb{K}(\varepsilon)\leq\overline{K}_{\varepsilon}\triangleq\left\lceil 2\log(C/\varepsilon)/\log(1/\overline{Q})\right\rceil.$ (3.3)
@@
By [Cortild:2024aa, Theorem 2.1], we also know that weak convergence of the iterates $(v_{k})$ and the velocities $(z_{k})$ is guaranteed when $\mathsf{T}_{k}$ is just nonexpansive. Moreover, if $\mathsf{T}_{k}$ is a fixed-point encoding of Problem (Aux), which has a unique solution $p$ by construction, we know that the sequence convergences weakly to this unique fixed-point. Specifically, in the finite-dimensional setting, we know that $\|v_{k}-p\|\to 0$ and $\|v_{k+1}-z_{k}\|\to 0$, meaning that Algorithm 2 still can be applied. However, Lemma 2.1 no longer holds, and we do not have an appropriate stopping criterion to guarantee $\|w^{+}-\overline{u}_{\beta}(w)\|\leq\mathtt{e}$. Within our final numerical implementation in Section 5.3, we experimentally test our algorithm on problems displaying only nonexpansive operators, corresponding to $q=1$, noting that the theoretical guarantees are weakened in this setting.
@@
Rather than assuming exact access to the operators $\mathsf{T}_{k}$, we could assume access to $\delta_{k}$-perturbations $\tilde{\mathsf{T}}_{k}$, with bounded inexactness model $0\leq\delta_{k}\leq\delta_{\max}$. Such an error structure is motivated by the fact that the exact map might in general not be available, or its computation may be very demanding. This occurs for instance when applying proximal methods to image deblurring with total variation [Chambolle:2004aa], or to structured sparsity regularization problems in machine learning and inverse problems [Zhao:2009aa]. In those cases, the proximity operator is usually computed using ad hoc algorithms, and therefore inexactly. The implications of inexact implementations are the following: Lemma 2.1 would still hold, by replacing $\varepsilon$ by $\max(\varepsilon,\delta_{\max})$, provided $\delta_{\max}$ is finite. Additionally, Theorem 3.1 would still hold [Cortild:2024aa], provided $(\theta_{k}\delta_{k})\in\ell^{2}$, but without an explicit rate. As such, Corollary 3.2 still holds, but again without rate guarantees. Specifically, we have no a-priori upper bound on the required number of iterations required for Algorithm 1 to terminate, but we know it will terminate.
@@ 3.2 Convergence Analysis of Restarting Procedure
For our complexity analysis, we introduce two gap functions, one associated with the solution set $\mathcal{S}_{0}$, and one with feasibility $\operatorname{Zer}(\mathsf{M})$.
Definition 3.4.
@@
$\operatorname{Gap}_{\operatorname{opt}}(u)\triangleq\sup_{v\in\mathcal{S}_{0}}\langle\mathsf{G}v,u-v\rangle.$ (3.4)
We observe that if $u$ is feasible, namely if $u\in\mathcal{S}_{0}=\operatorname{Zer}(\mathsf{M})$, then the condition $\operatorname{Gap}_{\operatorname{opt}}(u)\leq 0$ is equivalent to $u\in\operatorname{Zer}(\mathsf{G}+\operatorname{\mathsf{NC}}_{\mathcal{S}_{0}})$, such that $u$ solves Problem (P). It is important to note, however, that $\operatorname{Gap}_{\operatorname{opt}}(u)\leq 0$ provides limited information when $u\notin\operatorname{Zer}(\mathsf{M})$. It is thus important to obtain a lower bound on the optimality gap. The feasibility gap is motivated by the observation that, by maximal monotonicity of $\mathsf{M}$, for any $x\in\mathcal{H}$,
@@
Definition 3.5.
The feasibility gap $\operatorname{Gap}_{\operatorname{feas}}:\mathcal{H}\to\mathbb{R}$ is defined as
$\operatorname{Gap}_{\operatorname{feas}}(x)=\sup_{y\in\operatorname{dom}(\mathsf{M}),v\in\mathsf{M}y}\langle v,x-y\rangle.$ (3.5)
We note that $\operatorname{Gap}_{\operatorname{feas}}(x)\leq 0$ implies $x\in\operatorname{Zer}(\mathsf{M})$, and that $\operatorname{Gap}_{\operatorname{feas}}(x)\geq 0$ holds for all $x\in\mathcal{H}$ since, by assumption, the problem admits at least one solution; in particular, $\operatorname{Zer}(\mathsf{M})\neq\varnothing$. Since $\operatorname{Gap}_{\operatorname{opt}}$ and $\operatorname{Gap}_{\operatorname{feas}}$ are pointwise suprema of affine functions, both are weakly lower semicontinuous and convex.
@@
The notion of weak sharpness of zero sets is closely related to the classical concept of weak sharpness for variational inequalities [marcotte_weak_1998]. Geometric conditions ensuring the error-bound characterization (3.6) can be found in [Al-Homidan:2017aa]. The recent papers [samadi_improved_2025] and [Alvarez:2001aa] rely also heavily on this assumption. An important class of examples arises when $\mathsf{M}$ is single-valued and $\rho=1$, notably in monotone linear complementarity problems in finite dimensions under nondegeneracy conditions [burke_weak_1993, pang1997error].
Remark 3.4.
In the case where the lower-level problem is determined by a convex optimization problem, weak-sharpness implies an Hölderian errror bound, an assumption already imposed by [Cabot2005] in the context of hierarchical minimization. Specifically, let us assume that $\mathsf{M}=\partial f$ the subgradient of a proper convex and lower semi-continuous function $f:\mathcal{H}\to\mathbb{R}\cup\{+\infty\}$. Then,
@@
$D_{\mathsf{M}}=\operatorname{diam}(\operatorname{dom}(\mathsf{M})),\hskip 17.00024pt\hskip 17.00024pt$ $C_{\mathsf{M}}=\sup_{x\in\operatorname{dom}(\mathsf{M})}\sup_{v\in\mathsf{M}x}\|v\|,\qquad$ $C_{\mathsf{G}}=\sup_{x\in\operatorname{dom}(\mathsf{M})}\mathsf{G}(x).$
@@
The parameter sequences satisfy the following conditions:
@@
For any $(x,v)\in\operatorname{Gr}(\mathsf{M})$, it holds that
$\lambda_{n}\langle v,w_{n+1}-x\rangle+\lambda_{n}\beta_{n}\langle\mathsf{G}(x),w_{n+1}-x\rangle\leq-\frac{\alpha\lambda_{n+1}}{2}\|w_{n+1}-x\|^{2}+\frac{\alpha\lambda_{n}}{2}\|w_{n}-x\|^{2}+C_{1}\lambda_{n}\mathtt{e}_{n},$ (3.8)
@@
As $\beta_{n}\mathsf{G}(\overline{u}_{n+1})-\alpha(\overline{u}_{n+1}-w_{n})\in\mathsf{M}(\overline{u}_{n+1})$ and $v\in\mathsf{M}(x)$, it follows by monotonicity of $\mathsf{M}$ and by $\mu$-monotonicity of $\mathsf{G}$ that
$\langle w_{n}-\overline{u}_{n+1},x-\overline{u}_{n+1}\rangle$ $\leq\frac{1}{\alpha}\langle v,x-\overline{u}_{n+1}\rangle+\frac{\beta_{n}}{\alpha}\langle\mathsf{G}(\overline{u}_{n+1}),x-\overline{u}_{n+1}\rangle$
$\leq\frac{1}{\alpha}\langle v,x-\overline{u}_{n+1}\rangle+\frac{\beta_{n}}{\alpha}\langle\mathsf{G}(x),x-\overline{u}_{n+1}\rangle-\frac{\mu\beta_{n}}{\alpha}\|x-\overline{u}_{n+1}\|^{2}.$
@@
$\frac{1}{2}\|w_{n+1}-x\|^{2}-\frac{1}{2}\|w_{n}-x\|^{2}$ $\leq\frac{1}{\alpha}\langle v,x-\overline{u}_{n+1}\rangle+\frac{\beta_{n}}{\alpha}\langle\mathsf{G}(x),x-\overline{u}_{n+1}\rangle-\frac{\mu\beta_{n}}{\alpha}\|{x-w_{n+1}}\|^{2}$
@@
$\left(\frac{1}{2}+\frac{\mu\beta_{n}}{\alpha}\right)\|w_{n+1}-x\|^{2}-\frac{1}{2}\|w_{n}-x\|^{2}\leq\frac{1}{\alpha}\langle v,x-\overline{u}_{n+1}\rangle+\frac{\beta_{n}}{\alpha}\langle\mathsf{G}(x),x-\overline{u}_{n+1}\rangle+2\left(1+\frac{\mu\beta_{\max}}{\alpha}\right)(D_{\mathsf{M}}+\mathtt{e}_{\max})\mathtt{e}_{n}.$
@@
$\langle\mathsf{G}(x),x-\overline{u}_{n+1}\rangle=\langle\mathsf{G}(x),x-w_{n+1}\rangle+\langle\mathsf{G}(x),w_{n+1}-\overline{u}_{n+1}\rangle\leq\langle\mathsf{G}(x),x-w_{n+1}\rangle+C_{\mathsf{G}}\mathtt{e}_{n},$
@@
$\lambda_{n}\langle v,w_{n+1}-x\rangle$ $+\lambda_{n}\beta_{n}\langle\mathsf{G}(x),w_{n+1}-x\rangle$
@@
$\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}\langle\mathsf{G}(x),w_{n+1}-x\rangle$ $\leq\frac{-\alpha\lambda_{N}}{2}\|w_{N}-x\|^{2}+\frac{\alpha\lambda_{0}}{2}\|w_{0}-x\|^{2}+C_{1}\cdot\sum_{n=0}^{N-1}\lambda_{n}\mathtt{e}_{n}$
@@
$\langle\mathsf{G}(x),\overline{w}_{N}-x\rangle\leq C_{2}\cdot\frac{1}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}+C_{1}\cdot\frac{\sum_{n=0}^{N-1}\lambda_{n}\mathtt{e}_{n}}{\sum_{n=0}^{N-1}\lambda_{n}\beta_{n}}.$
@@
$\langle\mathsf{G}(x),w_{n+1}-x\rangle\geq-\|{\mathsf{G}(x)}\|\cdot\|{w_{n+1}-x}\|\geq-\|\mathsf{G}(x)\|\cdot(\|w_{n+1}-\overline{u}_{n+1}\|+\|\overline{u}_{n+1}-x\|)\geq-C_{\mathsf{G}}(\mathtt{e}_{n}+D_{\mathsf{M}}).$
It then follows $\langle\mathsf{G}(x),w_{n+1}-x\rangle\leq C_{\mathsf{G}}(\mathtt{e}_{\max}+D_{\mathsf{M}})$. Summing for $n=0,\ldots,N-1$ then yields
@@
We may now combine the two previous lemmas to state our main theorem of this paper. We establish a combined complexity result in terms of the two introduced gap functions using the averaged trajectory $\overline{w}_{N}$.
@@
    The upper bound is just Lemma 3.8. For the lower bound, we invoke the Cauchy-Schwarz inequality $\langle\mathsf{G}(x),\overline{w}_{N}-x\rangle\geq\|{\mathsf{G}(x)}\|\cdot\|{\overline{w}_{N}-x}\|$, implying
@@
Let $(w_{n})_{n=0}^{N}$ be generated through Algorithm 2. Assume that $\mathsf{G}$ is merely monotone (i.e. $\mu=0$). Let $\beta_{n}=(n+1)^{-b}$ for $b\in(0,1)$ and choose $\varepsilon_{n}$ such that $\mathtt{e}_{n}=o(\beta_{n})$. Then, for all $N\geq 2^{1/(1-b)}$, it holds that
@@
If, moreover, $\mathcal{S}_{0}$ is weakly sharp with constant $\kappa$ of order $\rho\geq 1$, then it also holds that $\operatorname{Gap}_{\operatorname{opt}}(\overline{w}_{N})\geq\mathcal{O}(\operatorname{Gap}_{\operatorname{feas}}(\overline{w}_{N})^{1/\rho}).$
@@
3.3 Total Complexity Analysis
We now finally perform the complete complexity analysis of our scheme DANTE by combining the bounds derived for Algorithms 1 and 2. Given a sequences of user-defined tolerances $(\varepsilon_{n})_{n=0}^{N}$, we known that Algorithm 1 terminates after at most $\overline{K}_{\varepsilon_{n}}\equiv\overline{K}_{n}$ iterations (Corollary 3.3, eq. (3.3)). To assess the total iteration complexity of Algorithm 2 for a pre-defined number of restarts $N\geq 1$, we thus need to compute the total iteration complexity
$\mathcal{C}(N,(\varepsilon_{n})_{n=0}^{N})\triangleq\sum_{n=0}^{N-1}\mathbb{K}(\varepsilon_{n}).$ (3.14)
To estimate the total iteration complexity, we make one final assumption.